# Generalization Error of the Tilted Empirical Risk

### v2

@@ 4 Generalization Bounds for Unbounded Loss Functions
Under the same assumption in Theorem 4.3 and a finite hypothesis space, then for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ with probability at least $(1-\delta)$ and $\gamma<0$, the excess risk of tilted empirical risk satisfies
$\mathfrak{E}_{\gamma}(\mu)$ $\leq\frac{4\kappa_{u}\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{B(\delta)}{n}}-\gamma\kappa_{u}^{2}-\frac{8\exp(-\gamma\kappa_{u})B(\delta)}{3n\gamma(1-\zeta)},$ (22)
@@
5 Robustness of TERM
As shown in experiments by Li et al. (2021), the tilted empirical risk is robust to noise or outliers samples during training using negative tilt ($\gamma<0$). In this section, we study the robustness of the TER under distributional shift, $\tilde{\mu}$ and the following assumption is made.
Assumption 5.1 (Uniform bounded second moment under $\tilde{\mu}$).
@@ 5 Robustness of TERM
There is a constant $\kappa_{s}\in\mathbb{R}^{+}$ such that the loss function $(H,Z)\mapsto\ell(H,Z)$ satisfies $\mathbb{E}_{\tilde{\mu}}[\ell^{2}(h,Z)]\leq\kappa_{s}^{2}$ uniformly for all $h\in\mathcal{H}$.
Using the functional derivative (Cardaliaguet et al., 2019), we can provide the following results.
Proposition 5.2.
Given Assumption 5.1 and Assumption 4.1, then the difference of tilted population risk under, (9), between $\mu$ and $\tilde{\mu}$ is bounded as follows,
$\frac{1}{|\gamma|}\Big{|}\log(\mathbb{E}_{\tilde{Z}\sim\mu}[C(h,\tilde{Z})])-\log(\mathbb{E}_{\tilde{Z}\sim\tilde{\mu}}[C(h,\tilde{Z})])\Big{|}\leq\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})}{(\kappa_{u}-\kappa_{s})},$ (23)
where $C(h,\tilde{Z})=\exp(\gamma\ell(h,\tilde{Z}))$.
Note that, for positive $\gamma$, the result in Proposition 5.2 does not hold and can be unbounded. Using Proposition 5.2, we can provide an upper and lower bounds on the tilted generalization error under distributional shift.
Proposition 5.3.
Given Assumptions  4.1 and 5.1, for any fixed $h\in\mathcal{H}$ and with probability least $(1-\delta)$ for $\gamma<0$, then the following upper bound holds on the tilted generalization error
$\begin{split}&\mathrm{gen}_{\gamma}(h,\hat{S})\leq 2\kappa_{s}\exp(-\gamma\kappa_{s})\sqrt{\frac{\log(2/\delta)}{n}}-\frac{\gamma}{2}\kappa_{u}^{2}\\ &\quad-\frac{4\exp(-\gamma\kappa_{s})(\log(2/\delta))}{3n\gamma}+\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}D(\kappa_{s},\kappa_{u}),\end{split}$
where $D(\kappa_{s},\kappa_{u})=\frac{\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})}{(\kappa_{u}-\kappa_{s})}$ and $\hat{S}$ is the training dataset under the distributional shift.
Proposition 5.4.
Given Assumptions  4.1 and 5.1, for any fixed $h\in\mathcal{H}$ and with probability least $(1-\delta)$, there exists a $\zeta\in(0,1)$ such that for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ and $\gamma<0$, such that the following upper bound holds on the tilted generalization error
$\begin{split}&\mathrm{gen}_{\gamma}(h,\hat{S})\geq-\frac{2\kappa_{s}\exp(-\gamma\kappa_{s})}{(1-\zeta)}\sqrt{\frac{\log(2/\delta)}{n}}\\ &\qquad+\frac{4\exp(-\gamma\kappa_{s})(\log(2/\delta))}{3n\gamma(1-\zeta)}-\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}D(\kappa_{s},\kappa_{u}),\end{split}$
where $D(\kappa_{s},\kappa_{u})=\frac{\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})}{(\kappa_{u}-\kappa_{s})}$ and $\hat{S}$ is the training dataset under the distributional shift.
@@
Combining Proposition 5.3 and Proposition 5.4, we derive an upper bound on the absolute value of the tilted generalization error under distributional shift.
@@
Under the same assumptions in Proposition 5.4, then for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ and $\gamma<0$, and with probability at least $(1-\delta)$, the absolute value of the tilted generalization error under distributional shift satisfies
@@
$\begin{split}&\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,\hat{S})|\leq\frac{2\kappa_{s}\exp(-\gamma\kappa_{s})}{(1-\zeta)}\sqrt{\frac{B(\delta)}{n}}-\frac{4\exp(-\gamma\kappa_{s})B(\delta)}{3n\gamma(1-\zeta)}\\ &\qquad-\frac{\gamma\kappa_{u}^{2}}{2}+\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})},\end{split}$ (24)
where $B(\delta)=\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)$.
Similar results to Proposition 5.3, can be derived via an information-theoretic approach.
Robustness vs Generalization: The term $\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})}$ represents the distributional shift cost (or robustness) associated with the TER. This cost can be reduced by increasing $|\gamma|$. However, increasing $|\gamma|$ also amplifies other terms in the upper bound of the tilted generalization error. Therefore, there is a trade-off between robustness and generalization, particularly for $\gamma<0$ in the TER. Interestingly, Li et al. (2021) also observed this trade-off for negative tilt.
6 The KL-Regularized TERM Problem
@@
Our upper bound in Corollary 3.15 on the absolute value of expected generalization error depends on the mutual information between $H$ and $S$. Therefore, it is of interest to investigate an algorithm which minimizes the regularized expected TERM via mutual information.
@@
where $\alpha$ is the inverse temperature. As discussed by Xu and Raginsky (2017); Aminian et al. (2023), the regularization problem in (25) is dependent on the data distribution, $P_{S}$. Therefore, we relax the problem in (25) by considering the following regularized version via KL divergence,
$P_{H|S}^{\star}=\mathop{\mathrm{arg}\,\inf}_{P_{H|S}}\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})+\frac{1}{\alpha}\mathrm{KL}(P_{H|S}\|\pi_{H}|P_{S}),$ (26)
where $I(H;S)\leq\mathrm{KL}(P_{H|S}\|\pi_{H}|P_{S})$ and $\pi_{H}$ is a prior distribution over hypothesis space $\mathcal{H}$. All proof details are deferred to Appendix G.
@@
is the solution to the KL-regularized ERM minimization problem, where $\tilde{F}_{\alpha}(S)$ is the normalization factor. Therefore, the tilted Gibbs posterior is different from the Gibbs posterior, (28). It can be shown that for $\gamma\rightarrow 0$, the tilted Gibbs posterior converges to the Gibbs posterior. Therefore, it is interesting to study the expected generalization error of the tilted Gibbs posterior. For this purpose, we give an exact characterization of the difference between the expected TER under the joint and the product of marginal distributions of $H$ and $S$. More results regarding the Gibbs posterior are provided in Appendix G.
@@
Under Assumption 3.1, the expected generalization error of the tilted Gibbs posterior satisfies
$\begin{split}&\overline{\mathrm{gen}}_{\gamma}(H,S)\leq\frac{\alpha(\exp(|\gamma|M)-1)^{2}}{2\gamma^{2}n}\\ &-\frac{\gamma\exp(-\gamma M)}{2}\big{(}1-\frac{1}{n}\big{)}\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\ell(H,\tilde{Z}))\big{]}.\end{split}$ (30)
Similar to Corollary 3.15, we derive the following upper bound on the absolute value of the expected tilted generalization error of the tilted generalization error.
Corollary 6.4.
Under the same assumptions in Theorem 6.3, the absolute value of the expected tilted generalization error of the tilted Gibbs posterior satisfies
$\begin{split}&|\overline{\mathrm{gen}}_{\gamma}(H,S)|\leq\frac{\alpha(\exp(|\gamma|M)-1)^{2}}{2\gamma^{2}n}\\ &+\frac{|\gamma|M^{2}\exp(|\gamma|M)}{8}\Big{(}1-\frac{1}{n}\Big{)}.\end{split}$ (31)
Remark 6.5 (Convergence rate).
Remark 6.6 (Discussion of $\gamma$).
From the upper bound in Theorem 6.3, we can observe that under $\gamma\rightarrow 0$ and Assumption 3.1, the upper bound converges to the upper bound on the Gibbs posterior (Aminian et al., 2021a). For positive tilt $(\gamma>0)$, and sufficient large value of $n$, the upper bound in Theorem 6.3, can be tighter than the upper bound on the Gibbs posterior.

### v3

@@ 4 Robustness of TER
Under the same assumptions in Theorem 3.5 for some $\epsilon\in(0,1]$, then for $n\geq\frac{(4|\gamma|^{1+\epsilon}\kappa_{u}^{1+\epsilon}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ and $\gamma<0$, and with probability at least $(1-\delta)$, the absolute value of the tilted generalization error under distributional shift satisfies
$\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,\hat{S})|$ (20)
$\leq\frac{2\exp(|\gamma|\kappa_{s})}{(1-\zeta)|\gamma|}\sqrt{\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{s}^{1+\epsilon}B(\delta)}{n}}+\frac{4\exp(|\gamma|\kappa_{s})B(\delta)}{3n|\gamma|(1-\zeta)}$
$+|\gamma|^{\epsilon}\kappa_{u}^{1+\epsilon}+\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})},$
@@
Using Lemma 3.7, we can derive an upper bound on excess risk under distribution shift.
Other Divergences: Our approach can be extended to incorporate other divergence measures that quantify distribution shift based on training and testing samples directly, rather than their underlying distributions. This is in line with methods commonly explored in the domain adaptation literature (e.g., (Ben-David et al., 2006; Zou et al., 2024; Ye et al., 2021)).
Comparison with ERM: Theorem 4.3 gives an upper bound on the tilted generalization error of TERM under distribution shifts, which is valid for all distribution shifts. Specifically, this upper bound depends on the total variation distance $\mathbb{TV}(\mu,\tilde{\mu})$, which is bounded for any distributions $\mu$ and $\tilde{\mu}$. This robustness is attributed to the negative tilt and properties of the exponential function. In contrast, for ERM, we need to derive an upper bound on the following term,
$\mathrm{Dif}(\mu,\tilde{\mu}):=\mathbb{E}_{Z\sim\mu}[\ell(h,Z)]-\mathbb{E}_{\tilde{Z}\sim\tilde{\mu}}[\ell(h,\tilde{Z})].$ (21)
For unbounded loss functions, deriving an upper bound on $\mathrm{Dif}(\mu,\tilde{\mu})$ in terms of total variation distance is not feasible. While such bounds can be established using KL-divergence $\mathrm{KL}(\mu\|\tilde{\mu})$ under specific conditions (e.g., when the loss function is sub-Gaussian under the data-generating distribution), these bounds may become unbounded for certain $\tilde{\mu}$. In this context, TERM with negative tilt emerges as a robust solution, providing theoretical performance guarantees for distribution-shift under unbounded loss functions.
Robustness vs Generalization: The term $\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}$ represents the distributional shift cost (or robustness) associated with the TER. This cost can be reduced by increasing $|\gamma|$. However, increasing $|\gamma|$ also amplifies other terms in the upper bound of the tilted generalization error. Therefore, there is a trade-off between robustness and generalization, particularly for $\gamma<0$ in the TER. Li et al. (2021) also empirically observed this trade-off for negative tilt.
Information-theoretic Approach: Using a similar approach, we can derive an upper bound on expected tilted generalization error under distribution shift. We make the following assumption.
Assumption 4.4.
@@
The learning algorithm $P_{H|\hat{S}}$, loss function $\ell$, and $\mu$ are such that there is a constant $\kappa_{t}\in\mathbb{R}^{+}$ with which the loss function $(H,Z)\mapsto\ell(H,Z)$ satisfies $\max\big{(}\mathbb{E}_{P_{H,\tilde{Z}}}[\ell^{1+\epsilon}(H,\tilde{Z})],\mathbb{E}_{P_{H}\otimes\tilde{\mu}}[\ell^{1+\epsilon}(H,\tilde{Z})]\big{)}\leq\kappa_{st}^{1+\epsilon}$ for some $\epsilon\in(0,1]$.
@@
Using Theorem 3.11 and similar results to Proposition 4.2, we derive an upper bound on absolute expected tilted generalization error under distribution shift.
@@
Under the same assumptions as in Theorem 3.11 and Assumption 4.4, the following upper bounds on absolute expected tilted generalization error hold under distributional shift, if (a) or (b) hold,
@@
$\begin{split}&|\overline{\mathrm{gen}}_{\gamma}(H,\hat{S})|\leq\frac{\exp(|\gamma|\kappa_{st})}{|\gamma|(1-\zeta)}G\big{(}I(H;\hat{S})\big{)}+|\gamma|^{\epsilon}\kappa\\ &+\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\Big{(}\exp(|\gamma|\kappa_{t})-\exp(|\gamma|\kappa_{st})\Big{)}}{(\kappa_{t}-\kappa_{st})},\end{split}$
where $\kappa=\kappa_{t}^{1+\epsilon}+\kappa_{st}^{1+\epsilon}$, if (a) or (b) hold we have $G(I(H;\hat{S}))=\sqrt{\frac{2\kappa_{st}^{\epsilon+1}|\gamma|^{1+\epsilon}I(H;\hat{S})}{n}}$ and if (c) holds, we have, $G(I(H;\hat{S}))=\frac{I(H;\hat{S})}{n}+\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{st}^{1+\epsilon}}{2}$.
We can observe the same robustness and generalization in Theorem 4.5.
5 Data-driven Choice of Tilt
In this section, we provide a data-driven approach for choosing the tilt ($\gamma$) based on theoretical results. It is noteworthy that the upper bound in Theorem 4.3 can be infinite for $\gamma\rightarrow-\infty$ and $\gamma=0$ and a fixed $n$. Consequently, there must exist a $\gamma\in(-\infty,0)$ that minimizes this upper bound. To illustrate this point, consider the case where $n\rightarrow\infty$; here, the first and second terms in the upper bound, (20), would vanish. Thus, we are led to the following minimization problem:
$\begin{split}&\gamma_{\text{data}}:=\mathop{\mathrm{arg}\min}_{\gamma\in(-\infty,0)}\Big{[}|\gamma|^{\epsilon}\kappa_{u}^{1+\epsilon}\\ &\qquad+\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})}\Big{]},\end{split}$ (22)
for which a solution $\gamma^{*}$ exists. As $\gamma^{\star}$ decreases when $\mathbb{TV}(\mu,\tilde{\mu})$ increases, practically, this implies that if the training distribution becomes more adversarial (i.e., further away from the benign test distribution), we would use smaller negative $\gamma$’s to bypass outliers. Therefore, we can consider $\gamma_{\text{data}}$ as a data-driven choice for the TERM problem. We consider a simple experiment to show the effectiveness of data-driven tilt, $\gamma_{\text{data}}$.
In this experiment, we consider the logistic regression setup described in Li et al. (2023a), adding a Gaussian or Pareto outlier dataset to the training dataset at different ratios ($\rho$) relative to the training dataset. We evaluate the following values:
  • $\gamma^{\star}$: The optimal tilt based on grid search
  • $\mathrm{R}(h_{\gamma^{\star}}(\hat{S}),\mu)$: The population risk under the optimal TERM solution where $h_{\gamma^{\star}}(\hat{S}):=\arg\min_{h\in\mathcal{H}}\hat{R}_{\gamma^{\star}}(h,\hat{S})$.
  • $\mathrm{R}(h_{\text{\tiny ERM}}(\hat{S}),\mu)$: The population risk under the ERM solution where $h_{\text{\tiny ERM}}(\hat{S}):=\arg\min_{h\in\mathcal{H}}\hat{R}(h,\hat{S})$.
  • $\gamma_{\text{data}}$: The data-driven tilt based on optimization of (22).
  • $\mathrm{R}(h_{\gamma_{\text{data}}}(\hat{S}),\mu)$: The population risk under the data-driven tilt solution where $h_{\gamma_{\text{data}}}(\hat{S}):=\arg\min_{h\in\mathcal{H}}\hat{R}_{\gamma_{\text{data}}}(h,\hat{S})$.
The training dataset consists of 1,000 samples. The results for Gaussian and Pareto outliers are reported in Table 1 and Table 2, respectively. For the Pareto-outlier as a distribution with unbounded second moment, we observe better performance for the data-driven approach in comparison with ERM. Note that the variance of TERM and data-driven TERM in comparison with ERM has less variance as expected. The details of the experiments are provided in Appendix G . More experiments for Linear regression as proposed in (Li et al., 2021) is provided in Appendix G.
Table 1: Logistic Regression with Gaussian Outliers: Results averaged over three runs ($n=1,000$ samples), showing mean $\pm$ standard deviation.
$\rho$ $\gamma^{\star}$ $\mathrm{R}(h_{\gamma^{\star}}(\hat{S}),\mu)$ $\mathrm{R}(h_{\text{\tiny ERM}}(\hat{S}),\mu)$ $\gamma_{\text{data}}$ $\mathrm{R}(h_{\gamma_{\text{data}}}(\hat{S}),\mu)$
$0.1\%$ $-0.53{\scriptstyle\pm 0.000}$ $0.00{\scriptstyle\pm 0.000}$ $0.05{\scriptstyle\pm 0.001}$ $-1.40{\scriptstyle\pm 0.000}$ $0.00{\scriptstyle\pm 0.000}$
$17.6\%$ $-2.98{\scriptstyle\pm 0.000}$ $0.15{\scriptstyle\pm 0.004}$ $0.22{\scriptstyle\pm 0.001}$ $-4.91{\scriptstyle\pm 0.002}$ $0.16{\scriptstyle\pm 0.003}$
$35.0\%$ $-3.86{\scriptstyle\pm 0.000}$ $0.16{\scriptstyle\pm 0.004}$ $0.30{\scriptstyle\pm 0.002}$ $-3.33{\scriptstyle\pm 0.000}$ $0.20{\scriptstyle\pm 0.002}$
$52.5\%$ $-2.10{\scriptstyle\pm 0.000}$ $0.11{\scriptstyle\pm 0.001}$ $0.28{\scriptstyle\pm 0.002}$ $-1.93{\scriptstyle\pm 0.000}$ $0.14{\scriptstyle\pm 0.001}$
$70.0\%$ $-1.23{\scriptstyle\pm 0.000}$ $0.14{\scriptstyle\pm 0.002}$ $0.18{\scriptstyle\pm 0.000}$ $-2.28{\scriptstyle\pm 0.000}$ $0.15{\scriptstyle\pm 0.002}$
Table 2: Logistic Regression with Pareto Outliers: Results averaged over three runs ($n=1,000$ samples), showing mean $\pm$ standard deviation.
$\rho$ $\gamma^{\star}$ $\mathrm{R}(h_{\gamma^{\star}}(\hat{S}),\mu)$ $\mathrm{R}(h_{\text{\tiny ERM}}(\hat{S}),\mu)$ $\gamma_{\text{data}}$ $\mathrm{R}(h_{\gamma_{\text{data}}}(\hat{S}),\mu)$
$0.1\%$ $-1.40{\scriptstyle\pm 0.000}$ $0.00{\scriptstyle\pm 0.000}$ $0.03{\scriptstyle\pm 0.001}$ $-0.70{\scriptstyle\pm 0.000}$ $0.01{\scriptstyle\pm 0.000}$
$17.58\%$ $-3.33{\scriptstyle\pm 0.000}$ $0.00{\scriptstyle\pm 0.003}$ $0.02{\scriptstyle\pm 0.000}$ $-0.88{\scriptstyle\pm 0.000}$ $0.01{\scriptstyle\pm 0.000}$
$35.05\%$ $-1.05{\scriptstyle\pm 0.000}$ $0.00{\scriptstyle\pm 0.000}$ $0.01{\scriptstyle\pm 0.002}$ $-0.70{\scriptstyle\pm 0.000}$ $0.01{\scriptstyle\pm 0.000}$
$52.53\%$ $-1.05{\scriptstyle\pm 0.000}$ $0.01{\scriptstyle\pm 0.000}$ $0.01{\scriptstyle\pm 0.002}$ $-1.06{\scriptstyle\pm 0.000}$ $0.01{\scriptstyle\pm 0.001}$
$70.00\%$ $-0.88{\scriptstyle\pm 0.000}$ $0.00{\scriptstyle\pm 0.002}$ $0.02{\scriptstyle\pm 0.001}$ $-0.70{\scriptstyle\pm 0.000}$ $0.01{\scriptstyle\pm 0.000}$
6 The KL-Regularized TERM Problem for Unbounded Loss Functions
@@
Our upper bound in Theorem 3.11 on the absolute value of the expected tilted generalization error depends on the mutual information between $H$ and $S$. Therefore, it is of interest to investigate an algorithm that minimizes the regularized expected TERM via mutual information.
@@
where $\alpha$ is the inverse temperature. As discussed by Xu & Raginsky (2017); Aminian et al. (2023), the regularization problem in (23) is dependent on the data distribution, $P_{S}$. Therefore, we relax the problem in (23) by considering the following regularized version via KL divergence,
$P_{H|S}^{\star}$ $=\mathop{\mathrm{arg}\,\inf}_{P_{H|S}}\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})$ (24)
$\quad+\frac{1}{\alpha}\mathrm{KL}(P_{H|S}\|\pi_{H}|P_{S}),$
where $I(H;S)\leq\mathrm{KL}(P_{H|S}\|\pi_{H}|P_{S})$ and $\pi_{H}$ is a prior distribution over hypothesis space $\mathcal{H}$. All proof details are deferred to Appendix F.
@@
is the solution to the KL-regularized ERM minimization problem, where $\tilde{F}_{\alpha}(S)$ is the normalization factor. Therefore, the tilted Gibbs posterior is different from the Gibbs posterior, (26). It can be shown that for $\gamma\rightarrow 0$, the tilted Gibbs posterior converges to the Gibbs posterior. Therefore, it is interesting to study the expected tilted generalization error of the tilted Gibbs posterior. For this purpose, we give an exact characterization of the difference between the expected TER under the joint and the product of marginal distributions of $H$ and $S$.
@@
Under the same Assumptions, cases (a) and (b) in Theorem 3.11, the expected tilted generalization error of the tilted Gibbs posterior satisfies
$\begin{split}&0\leq\overline{\mathrm{gen}}_{\gamma}(H,S)\leq\frac{2\alpha\exp(2|\gamma|\kappa_{t})\kappa_{t}^{1+\epsilon}}{(1-\zeta)^{2}n|\gamma|^{1-\epsilon}}\\ &\quad+\frac{\exp(|\gamma|\kappa_{t})\kappa_{t}^{1+\epsilon}|\gamma|^{1/2+\epsilon}}{(1-\zeta)|\gamma|}\sqrt{\frac{2\alpha}{n}}+2|\gamma|^{\epsilon}\kappa_{t}^{1+\epsilon}.\end{split}$ (28)
Convergence rate: If $\gamma=O(1/n)$, then we obtain a convergence rate of $O(n^{-\epsilon})$ for the upper bound on the tilted generalization error of the tilted Gibbs posterior.
Comparison with the Gibbs posterior: Our results offer several advantages over the prior work in Aminian et al. (2021a). While Theorem 3 in their work establishes an $O(1/n)$ upper bound on the expected tilted generalization error of the Gibbs posterior, it requires the strong assumption of sub-Gaussian loss functions. In contrast, we achieve the same $O(1/n)$ convergence rate under the weaker condition of bounded second moments (when $\epsilon=1$) for the tilted Gibbs posterior. Additionally, our analysis extends to loss functions with bounded $(1+\epsilon)$-th moments for some $\epsilon\in(0,1]$, a scenario not addressed in the Gibbs posterior framework.